# Probing Primordial black holes with the distortion of Stochastic Gravitational Wave Background

### v1

@@ I Introduction
The stochastic gravitational-wave background (SGWB) arises from the incoherent superposition of a large number of independent gravitational-wave (GW) sources and can have both cosmological [9] and astrophysical [10] origins. For a cosmological SGWB, the signal may be generated by quantum vacuum fluctuations in the early Universe, while for an astrophysical SGWB, the cumulative emission from compact binary coalescences, such as binary black hole (BBH) mergers, can contribute significantly. According to Ref. [11], the BBH-generated SGWB is expected to dominate within the LIGO–Virgo frequency band. The astrophysical SGWB is of particular interest for at least two reasons [12]. First, the spectral shape encodes global statistical information about the underlying source population, allowing one to infer properties such as merger rates and mass distributions. Second, it may serve as a foreground that obscures the detection of the cosmological SGWB originating from the early Universe.
In addition to GWs themselves, gravitational lensing is another fundamental prediction of general relativity, describing the deflection of light or radiation as it passes near massive objects. This phenomenon has been widely exploited in astrophysics to probe dark matter distributions, measure cosmological parameters, and test general relativity. Importantly, gravitational lensing can also affect GW signals [13], introducing characteristic distortions that depend on the mass and configuration of the lens. Two distinct regimes are generally considered in GW lensing. When the GW wavelength is much shorter than the Schwarzschild radius associated with the lens, the system can be well approximated in the geometric-optics limit, similar to electromagnetic lensing by galaxies or clusters. In contrast, when the GW wavelength is comparable to or exceeds the characteristic lens scale, diffraction effects become non-negligible, and the wave-optics framework must be adopted [14]. The latter regime can introduce frequency-dependent modulations to the waveform, a phenomenon that has been extensively explored in the literature [15, 16, 17, 18].
@@ II.1 Specific form of BBH merge rate
where $\lambda$ denotes the mass fraction of the stars that end up in compact binary systems (with unit $M_{\odot}^{-1}$) and $\dot{\rho}_{*}(z)$ represents the cosmic star formation rate density (in units of $M_{\odot}\,\mathrm{yr^{-1}\,Mpc^{-3}}$). In practice, the value of $\lambda$ varies depending on the detector sensitivity and assumed population models.
@@ II.2 Energy spectrum of a single BBH
$f=\frac{a\eta^{2}+b\eta+c}{\pi M},$ (10)
where the coefficients $a$, $b$, and $c$ are provided in Table 1 of Ajith et al. [32].
@@ IV Mechanisms of Gravitational Lensing on the SGWB
Now we have all the tools we need to build analytical form of lensed SGWB:
$\Omega_{\mathrm{GW}}^{\mathrm{L}}(f)=\frac{f}{\rho_{c,0}H_{0}}\Bigg[$ $\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\frac{R_{\mathrm{V}}(z_{s})}{(1+z_{s})\,E(z_{s})}\,\tau(z_{s})\left.\left(\frac{dE_{\mathrm{GW}}}{df_{r}}\right)\right|_{f_{r}=(1+z_{s})f}\left.\left\langle\left|F(f_{l})\right|^{2}\right\rangle\right|_{f_{l}=(1+z_{l}^{*})f}\,dz_{s}$
$+\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\frac{R_{\mathrm{V}}(z_{s})}{(1+z_{s})\,E(z_{s})}\,(1-\tau(z_{s}))\left.\left(\frac{dE_{\mathrm{GW}}}{df_{r}}\right)\right|_{f_{r}=(1+z_{s})f}\,dz_{s}\Bigg].$ (21)
In realistic scenarios, each individual gravitational wave event, if lensed, would be affected by a lens located at a different redshift and characterized by distinct lens parameters such as mass and position. However, in order to derive analytical expressions, we adopt a simplifying assumption that all events are lensed by identical lenses, with fixed redshift and mass (and other relevant parameters). As a result, the analytical expressions presented above do not incorporate any distribution over lens properties. Naturally, future studies may relax this assumption by introducing lens parameter distributions to improve the accuracy of the model.
@@
According to Takahashi & Nakamura (2003) [16], the optical depth $\tau$ can exceed unity, implying that a single event may be lensed multiple times. Consequently, Eq. IV needs to be modified to account for this multiplicity:
$\Omega_{\mathrm{GW}}^{\mathrm{L}}(f)=\frac{f}{\rho_{c,0}H_{0}}\Bigg\{$ $\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\frac{R_{\mathrm{V}}(z_{s})}{(1+z_{s})E(z_{s})}\left[1-\lceil\tau(z_{s})\rceil+\tau(z_{s})\right]\left.\left(\frac{dE_{\mathrm{GW}}}{df_{r}}\right)\right|_{f_{r}=(1+z_{s})f}\left[\left\langle\left|F(f_{l})\right|^{2}\right\rangle\Big|_{f_{l}=(1+z_{l}^{*})f}\right]^{\lceil\tau(z_{s})\rceil}\,dz_{s}$
$+\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\frac{R_{\mathrm{V}}(z_{s})}{(1+z_{s})E(z_{s})}\left[\lceil\tau(z_{s})\rceil-\tau(z_{E}qs)\right]\left.\left(\frac{dE_{\mathrm{GW}}}{df_{r}}\right)\right|_{f_{r}=(1+z_{s})f}\left[\left\langle\left|F(f_{l})\right|^{2}\right\rangle\Big|_{f_{l}=(1+z_{l}^{*})f}\right]^{\lfloor\tau(z_{s})\rfloor}\,dz_{s}\Bigg\},$ (25)
where $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the ceiling and floor functions, respectively.
@@
where we set $y_{\mathrm{cr}}=3$.
@@
$F(f)=\exp$ $\left[\frac{\pi w}{4}+i\frac{w}{2}\left(\ln\left(\frac{w}{2}\right)-2\phi_{m}(y)\right)\right]$ (29)
$\times\Gamma\left(1-\frac{i}{2}w\right)\,{}_{1}F_{1}\left(\frac{i}{2}w,1;\frac{i}{2}wy^{2}\right),$

### v2

@@ I Introduction
The stochastic gravitational-wave background (SGWB) arises from the incoherent superposition of a large number of independent gravitational-wave (GW) sources and can have both cosmological [12] and astrophysical [25, 34] origins. For a cosmological SGWB, the signal may be generated by quantum vacuum fluctuations in the early Universe, while for an astrophysical SGWB, the cumulative emission from compact binary coalescences, such as binary black hole (BBH) mergers, can contribute significantly. According to Ref. [1], the BBH-generated SGWB is expected to dominate within the LIGO–Virgo frequency band. The astrophysical SGWB is of particular interest for at least two reasons [37]. First, the spectral shape encodes global statistical information about the underlying source population, allowing one to infer properties such as merger rates and mass distributions. Second, it may serve as a foreground that obscures the detection of the cosmological SGWB originating from the early Universe.
In addition to GWs themselves, gravitational lensing is another fundamental prediction of general relativity, describing the deflection of light or radiation as it passes near massive objects. This phenomenon has been widely exploited in astrophysics to probe dark matter distributions, measure cosmological parameters, and test general relativity. Importantly, gravitational lensing can also affect GW signals [22, 18, 19], introducing characteristic distortions that depend on the mass and configuration of the lens. Two distinct regimes are generally considered in GW lensing. When the GW wavelength is much shorter than the Schwarzschild radius associated with the lens, the system can be well approximated in the geometric-optics limit, similar to electromagnetic lensing by galaxies or clusters. In contrast, when the GW wavelength is comparable to or exceeds the characteristic lens scale, diffraction effects become non-negligible, and the wave-optics framework must be adopted [7]. The latter regime can introduce frequency-dependent modulations to the waveform, a phenomenon that has been extensively explored in the literature [21, 27, 28, 17].
@@ II.1 Specific form of BBH merge rate
where $\lambda$ denotes the mass fraction of the stars that end up in compact binary systems (with unit $M_{\odot}^{-1}$) and $\dot{\rho}_{*}(z)$ represents the cosmic star formation rate density (SFR)(in units of $M_{\odot}\,\mathrm{yr^{-1}\,Mpc^{-3}}$). In practice, the value of $\lambda$ varies depending on the detector sensitivity and assumed population models.
@@ II.2 Energy spectrum of a single BBH
$f=\frac{a_{k}\eta^{2}+b_{k}\eta+c_{k}}{\pi M},$ (10)
where the coefficients $a_{k}$, $b_{k}$, and $c_{k}$ are provided in Table 1.
Table 1: Coefficients $a_{k}$, $b_{k}$, and $c_{k}$ for the fitting formula 10[4].
Parameter $a_{k}$ $b_{k}$ $c_{k}$
$f_{\mathrm{1}}$ $2.9740\times 10^{-1}$ $4.4810\times 10^{-2}$ $9.5560\times 10^{-2}$
$f_{\mathrm{2}}$ $5.9411\times 10^{-1}$ $8.9794\times 10^{-2}$ $1.9111\times 10^{-1}$
$\sigma$ $5.0801\times 10^{-1}$ $7.7515\times 10^{-2}$ $2.2369\times 10^{-2}$
$f_{\mathrm{3}}$ $8.4845\times 10^{-1}$ $1.2848\times 10^{-1}$ $2.7299\times 10^{-1}$
@@ IV Mechanisms of Gravitational Lensing on the SGWB
Now we have all the tools we need to build analytical form of lensed SGWB: {align} Ω_GW^L(f) = fρc,0H0 [ ∫_z_min^z_max RV(zs)(1+zs) E(zs) τ(z_s) ( dEGWdfr )—_f_r=(1+z_s)f ⟨—F(f_l)—^2 ⟩—_f_l=(1+z_l^*)f dz_s
+ ∫_z_min^z_max RV(zs)(1+zs) E(zs) (1 - τ(z_s)) ( dEGWdfr )—_f_r=(1+z_s)f dz_s ]. In realistic scenarios, each individual gravitational wave event, if lensed, would be affected by a lens located at a different redshift and characterized by distinct lens parameters such as mass and position. However, in order to derive analytical expressions, we adopt a simplifying assumption that all events are lensed by identical lenses, with fixed redshift and mass (and other relevant parameters). As a result, the analytical expressions presented above do not incorporate any distribution over lens properties. Naturally, future studies may relax this assumption by introducing lens parameter distributions to improve the accuracy of the model.
@@
According to Takahashi & Nakamura (2003) [27], the optical depth $\tau$ can exceed unity, implying that a single event may be lensed multiple times. Consequently, Eq. 2 needs to be modified to account for this multiplicity: {align} Ω_GW^L(f) = fρc,0H0 { ∫_z_min^z_max RV(zs)(1+zs) E(zs) [1 - ⌈τ(z_s) ⌉+ τ(z_s)] ( dEGWdfr )—_f_r = (1 + z_s)f [⟨—F(f_l)—^2 ⟩_f_l = (1 + z_l^*)f ]^⌈τ(z_s) ⌉ dz_s
+ ∫_z_min^z_max RV(zs)(1+zs) E(zs) [⌈τ(z_s) ⌉- τ(z_s)] ( dEGWdfr )—_f_r = (1 + z_s)f [⟨—F(f_l)—^2 ⟩_f_l = (1 + z_l^*)f ]^⌊τ(z_s) ⌋ dz_s }, where $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the ceiling and floor functions, respectively.
@@
where we set $y_{\mathrm{cr}}=3$. Astrophysical black holes may also contribute to the lensed SGWB, $\Omega^{\mathrm{L}}_{\mathrm{GW}}$. In this work, however, we focus on the upper-limit impact of compact-object lensing and therefore neglect the astrophysical black-hole contribution. The PBH abundance as a dark matter component is still uncertain, but for a not very small PBH fraction $f_{\mathrm{PBH}}$, the lensing probability can be substantially larger than that of astrophysical black holes. Even under the extreme assumption that all baryonic matter forms black holes, their total abundance would be only $\sim 0.04$, compared to a possible PBH fraction of $\sim 0.26$ (assuming total matter density $\Omega_{M}=0.3$). In the future, one may distinguish them with the knowledge of stars or according to their mass distributions.
@@
$aligned F(f)=\exp&\left[\frac{\pi w}{4}+i\frac{w}{2}\left(\ln\left(\frac{w}{2}\right)-2\phi_{m}(y)\right)\right]\\ &\times\Gamma\left(1-\frac{i}{2}w\right)\,{}_{1}F_{1}\left(\frac{i}{2}w,1;\frac{i}{2}wy^{2}\right),$ (27)